OpenStudy (anonymous):

What is the general form of the equation of a circle with a center at the (4, 2) and a radius of 9? (x – 4)2 + (y – 2)2 = 81 (x – 4)2 + (y + 2)2 = 81 (x + 4)2 + (y – 2)2 = 81 (x + 4)2 + (y + 2)2 = 81

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

I'm guessing its the first one?

OpenStudy (anonymous):

No! Actually the last :) since 4 and 2 are positives

OpenStudy (anonymous):

x=4 x-4=0